# Comparison of chiral limit studies in curvature mass versus on-shell renormalized quark-meson model using ChPT

### v1

@@ III.3 Comparing various aspects of Columbia plots in the QMVT and RQM model for different $\bf{m_{\sigma}}$
The Fig. (7(a)) depicts, the comparison of the lines and regions in the horizontal $m_{\pi}-m_{K}$ plane (at $\mu=0$) of the three dimensional Columbia plots of the QMVT (RQM) model in the Fig. (4) (Fig. (4)) for the $m_{\sigma}=400$ MeV and Fig. (5) (Fig. (5)) for the $m_{\sigma}=530$ MeV. The Z(2) critical black solid line, in the horizontal $\mu-m_{K}$ plane of the above figures, has been replotted, by the dash dotted deep green line (solid red line ) for the $m_{\sigma}=530(400)$ MeV in the QMVT model and the dash double dotted sky blue line (solid indigo line) for the $m_{\sigma}=(530)400$ MeV in the RQM model. It is recalled that the QMVT and RQM model Z(2) critical lines for the $m_{\sigma}=400(530)$ MeV, intersect the SU(3) symmetric $m_{\pi}=m_{K}$ chiral limit line at the critical pion mass $m_{\pi}^{c}$ and terminate on the strange chiral limit line satisfying the $h_{y}=0=h_{y0}$, at the terminal pion mass $m_{\pi}^{t}$. One finds the $(m_{\pi}^{t},m_{\pi}^{c})=(98.15,78.45)$($(m_{\pi}^{t},m_{\pi}^{c})=(90.03,72.12)$) MeV in the QMVT model while in the RQM model one gets the $(m_{\pi}^{t},m_{\pi}^{c})=(168.39,134.16)$ ($(m_{\pi}^{t},m_{\pi}^{c})=(160.75,128.38)$) MeV for the $m_{\sigma}=400(530)$ MeV. The Fig. (7(a)) clearly demonstrates that the first order area underneath the crossover region of the QMVT model, is significantly smaller than the corresponding first order area lying below the crossover region of the RQM model. Table (3) shows the comparison of our present QMVT model results with those of the e-MFA:QM model FRG work Resch and RQM model work reported very recently in the Ref. vkt25 . The FRG work reports the $(m_{K}^{TCP},~{}m_{\pi}^{t},~{}m_{\pi}^{c})=(169,~{}110,~{}86)$ MeV for the $m_{\sigma}=530$ MeV, and we are finding $(m_{K}^{TCP},~{}m_{\pi}^{t},~{}m_{\pi}^{c})=(124.7,~{}90.03,~{}72.12)$ MeV in the present QMVT model work while one notes the $(m_{K}^{TCP},~{}m_{\pi}^{t},~{}m_{\pi}^{c})\equiv(235.7,~{}160.75,~{}128.38)$ MeV in the RQM model work. While being much smaller than their numerical values computed in the RQM model, the critical quantities $(m_{K}^{TCP},~{}m_{\pi}^{t},~{}m_{\pi}^{c})$ found in the present QMVT model work, are somewhat smaller but stand closer to the corresponding values reported in the e-MFA QM model FRG study of the Ref.Resch . The tricritical lines in the $\mu-m_{K}$ plane, show similar saturation trends for the QMVT model and the e-MFA:QM model FRG study while the saturation trend for the RQM model tricritical line gets spoiled.
@@
The first order regions in the $\mu-m_{K}$ and $m_{\pi}-m_{K}$ planes of the QMVT model Columbia plot,  while being much smaller than the corresponding first order regions for the RQM model, are similar but moderately smaller than the extent of the first order regions computed for the e-MFA:QM model FRG study. Here, it is relevant to point out that the $\sigma$ mass, in the e-MFA:QM model FRG study, is calculated from the curvature of the effective potential that gets evaluated by the FRG flow equation. Recall that the curvature masses of mesons are used for fixing the parameters in the QMVT model and this context of incorporating quark one-loop vacuum correction in the QMVT model is analogous to that of the e-MFA:QM model FRG study in the Ref.Resch . It is important to point out that even though the fraction $\beta_{cep}$, by which the pion and kaon masses are to be reduced in the $m_{\pi}-m_{K}$ plane such that the chiral crossover ends into a second order critical end point giving the first order transition region afterwards when $\beta<\beta_{cep}$, is almost the same in both the models as $(\beta_{cep}|_{m_{\sigma}=530\text{MeV}:\text{QMVT}}=0.2028)\simeq(\beta_{cep}|_{m_{\sigma}=530\text{MeV}:\text{eMFA;QM-FRG}}\equiv 0.20)$, the extent first order transition regions in the QMVT model Columbia plot, is somewhat smaller than what is found in the e-MFA:QM model Columbia plot as the critical quantities $(m_{K}^{TCP},~{}m_{\pi}^{t},~{}m_{\pi}^{c})|_{\text{QMVT}}=(124.7,~{}90.03,~{}72.12)<(m_{K}^{TCP},~{}m_{\pi}^{t},~{}m_{\pi}^{c})|_{\text{eMFA:QM-FRG}}=(169,~{}110,~{}86)$ MeV for the $m_{\sigma}=530$ MeV. The possible reason of the above difference lies in the different methods adopted for performing chiral limit study. Recall that the $\sigma$ curvature mass $m_{\sigma}=530$ MeV does not change towards the chiral limit in the QMVT model when the ($m_{\pi},m_{K}$) dependent ChPT scaling relations for the $f_{\pi},f_{K}\text{ and }M_{\eta}$, are used for calculating the change in model parameters away from the physical point whereas in contrast, the vacuum value of the $m_{\sigma}$ decreases significantly towards the chiral limit in the e-MFA:QM model work where the $f_{\pi},M_{\eta}$ and the model parameters are kept fixed as at the physical point when the scale $\Lambda$ of the initial effective action for the FRG flow is successively increased to Λ=700,1000,1100 and $\Lambda=700,1000,1100\text{ and }1143$ for compensating the successive decrease in the chiral symmetry breaking source strengths $j_{x}\text{ and }j_{y}$ by the fractions $\alpha=\sqrt{\beta}=1.0,0.17,0.04\text{ and }0$. The quark one-loop vacuum fluctuation generates quite a large smoothing effect on the strength of chiral transition in the e-MFA:QM model FRG study and this softening of the chiral transition by the fluctuation is very large (larger than that of the e-MFA:QM FRG study) in the QMVT model. Note that the full FRG QM model study where meson thermal and vacuum quantum fluctuations are also included, finds a very small $m_{\pi}^{c}\equiv 17$ MeV for which Pisarski et. al. pisarski24 has observed that this is partly because the approximation used in Ref. Resch is known to overestimate the masonic fluctuations that tend to soften the chiral transition PawlRen . Note also that the Ref. fejoHastuda has cautioned that the LPA in the FRG studies, completely neglects the wave function renormalization which might affect the final result. Recall that the quark one-loop vacuum fluctuation softens the chiral transition moderately in the RQM model because here the on-shell renormalization of parameters gives rise to stronger $U_{A}(1)$ anomaly strength $c$ together with a weaker light and strange chiral symmetry breaking strengths $h_{x}\text{ and }h_{y}$. Hence one gets significantly large first order regions in the RQM model Columbia plots vkt25 .
@@ IV Summary
The consistent chiral limit has been explored in the QMVT model schafwag12 ; chatmoh1 ; vkkt13 ; Rai whose parameters get fixed by the use of curvature masses of mesons after the quark one-loop vacuum correction is included in its effective potential in the $\overline{MS}$ scheme using the extended mean field approximation (e-MFA). In the RQM model treatment of quark one-loop vacuum fluctuation, one finds the renormalized parameters by matching the counterterms in the $\overline{MS}$ scheme with those of the on-shell scheme vkkr22 ; skrvkt24 . Using the $\mathcal{O}(\frac{1}{f^{2}})$ accurate results of the infrared regularized U(3) ChPT Herpay:05 ; borasoyI ; borasoyII ; Beisert ; Becher scaling relations for the ($m_{\pi},m_{K}$) dependent expressions of the $f_{\pi},f_{K}\text{ and }M_{\eta}$, the parameter fixing for the QMVT and RQM models, have been rendered free from any ambiguity or heuristic adjustment as the $\pi,K$ meson masses are reduced to move away from the physical point for performing the chiral limit study. After generating the Columbia plots in the QMVT model for the $m_{\sigma}=400\text{ and }530$ MeV, these plots have been compared and contrasted with the Columbia plots generated respectively in the e-MFA:QM model FRG study Resch for the $m_{\sigma}=530$ MeV and the RQM model study for the $m_{\sigma}=400\text{ and }530$ MeV as reported very recently in the Ref.vkt25 . Different methods of treating the fermionic vacuum fluctuations, cause significant differences in the extent of first order, second order and crossover regions that one finds in the Columbia plot.
@@
Very close to the chiral limit when $\beta=0.00725$, one gets huge first order gap in the QMVT and RQM model order parameters. Note that for the smaller $\beta<\beta_{cep}$, the difference in the QMVT and RQM model $T_{c}^{\chi}$ decreases from about 24 MeV for the $\beta=1.0\text{ and }0.5$ to 16.35 MeV for $\beta=0.15$ and 13.35 MeV for $\beta=0.00725$. In the exact chiral limit $f_{\pi}=f_{K}$, the $m_{\eta}=0$ and the heavy $m_{\eta^{\prime}}\equiv 825$ MeV signifies the $U_{A}(1)$ anomaly for both the QMVT model and RQM model. The spontaneous chiral symmetry breaking ( SCSB) for the $m_{\sigma}=400-800$ MeV, does not get lost in the QMVT and RQM model when the infrared regularized U(3) ChPT scaling Herpay:05 of the $m_{\pi},m_{K},M_{\eta}^{2}$ is used for the $(m_{\pi},m_{K}\rightarrow 0)$ chiral limit study. We point out that the $\sigma$ meson vacuum curvature mass stays the same as $m_{\sigma}=530$ MeV when the fraction $\beta$ is reduced in the QMVT model, whereas in the RQM model, the $\sigma$ curvature mass $m_{\sigma,c}$  is lowest at the physical point and it increases successively when one moves towards the chiral limit by reducing $\beta$. In complete contrast of the above two scenarios, the $\sigma$ mass in the e-MFA:QM model FRG study, registers a significant successive decrease towards the chiral limit from its physical point value of $m_{\sigma}=530$ MeV. The QMVT and RQM model $m_{\sigma}$ becomes zero in its temperature variation at the $\beta_{cep}$ of the corresponding model.
@@
The second order transition turns first order at the tricritical point $m_{K}=m_{K}^{TCP}$ ($m_{s}=m_{s}^{TCP}$) on the $m_{K}(m_{s})$ axis of the Columbia plot. The line of $Z(2)$ critical points, intersects the SU(3) symmetric $m_{\pi}=m_{K}$ ($m_{ud}=m_{s}$) line and terminates on the strange chiral limit line $h_{y0}=0(m_{s}=0)$ respectively at the $m_{\pi}^{c}$ ($m_{ud}^{c}$) and $m_{\pi}^{t}$ ($m_{ud}^{t}$) known as critical and terminal $\pi(\text{light quark})$ mass. In the $m_{\pi}-m_{K}$ and $\mu-m_{K}$ plane of the QMVT model Columbia plot for the $m_{\sigma}=400$ MeV, the computed critical quantities in the present work are $(\ m_{K}^{TCP},m_{\pi}^{t},m_{\pi}^{c})=(137.2,98.15,78.45)$ MeV which give the corresponding critical quantities in the $m_{ud}-m_{s}$ and $\mu-m_{s}$ plane as $(m_{s}^{TCP},m_{ud}^{t},m_{ud}^{c})=(8.06,2.05,1.31)$ MeV. In the RQM model $m_{\pi}-m_{K}$ and $\mu-m_{K}$ plane of the Columbia plot when the $m_{\sigma}=400$ MeV, one finds $(\ m_{K}^{TCP},m_{\pi}^{t},m_{\pi}^{c})=(242.6,168.39,134.16)$ MeV which give the $(m_{s}^{TCP},m_{ud}^{t},m_{ud}^{c})=(25.53,6.13,3.84)$ MeV in the $m_{ud}-m_{s}$ and $\mu-m_{s}$ plane. For the $m_{\sigma}=530$ MeV in the RQM model, the critical quantities reported in the Ref. vkt25 , are $(\ m_{K}^{TCP},m_{\pi}^{t},m_{\pi}^{c})=(235.7,160.75,128.38)$ MeV which correspond to the light-strange quark mass plane critical quantities as $(m_{s}^{TCP},m_{ud}^{t},m_{ud}^{c})=(24.08,5.58,3.51)$ MeV. In the present QMVT model Columbia plot calculation for the $m_{\sigma}=530$ MeV, we are finding the $(\ m_{K}^{TCP},m_{\pi}^{t},m_{\pi}^{c})=(124.7,90.03,72.12)$ MeV which give the $m_{ud}-m_{s}$ and $\mu-m_{s}$ plane critical quantities as $(m_{s}^{TCP},m_{ud}^{t},m_{ud}^{c})=(6.64,1.72,1.1)$ MeV. Note that the critical light quark mass $m_{ud}^{c}=1.1-1.31$ MeV is small in the QMVT model and it is $m_{ud}^{c}=3.51-3.84$ MeV in the RQM model whereas the recent LQCD study in Ref. zhang24 using Möbius domain wall fermions finds that $m_{ud}^{c}\leq$4 MeV. The comparison of the critical quantities and the $\beta_{cep}$ of the QMVT model with those of the RQM model, clearly establish that the first order regions in the $\mu-m_{K}$($\mu-m_{s}$) and $m_{\pi}-m_{K}$($m_{ud}-m_{s}$) planes of the QMVT model Columbia plot, are much smaller than the corresponding RQM model first order regions which are significantly large because the smoothing effect of the quark one-loop vacuum fluctuation on the strength of chiral transition is quite moderate as demonstrated in the Ref. vkt25 .
@@
Comparing the results for the QMVT model and the e-MFA:QM FRG model, we note that even though the fraction $\beta_{cep}$ is almost the same in both the models as $(\beta_{cep}|_{m_{\sigma}=530\text{MeV}:\text{QMVT}}=0.2028)\simeq(\beta_{cep}|_{m_{\sigma}=530\text{MeV}:\text{eMFA;QM-FRG}}\equiv 0.20)$, the extent of first order transition regions in the QMVT model Columbia plot, is somewhat smaller than what is found in the e-MFA:QM model FRG Columbia plot as the critical quantities $(m_{K}^{TCP},~{}m_{\pi}^{t},~{}m_{\pi}^{c})|_{\text{QMVT}}=(124.7,~{}90.03,~{}72.12)<(m_{K}^{TCP},~{}m_{\pi}^{t},~{}m_{\pi}^{c})|_{\text{eMFA:QM-FRG}}=(169,~{}110,~{}86)$ MeV for the $m_{\sigma}=530$ MeV. The possible reason of the above difference lies in the different methods adopted for performing the chiral limit study. Recall that the $m_{\sigma}=530$ MeV does not change in the QMVT model when the ChPT scaling relations for the $f_{\pi},f_{K}\text{ and }M_{\eta}$ are used to approach the chiral limit whereas in contrast, the vacuum $m_{\sigma}$ decreases significantly towards the chiral limit in the e-MFA:QM-FRG work when the scale $\Lambda$ of the initial effective action for the FRG flow is successively increased. While being nearly similar and comparable in the e-MFA:QM-FRG and the QMVT model study, the significantly large smoothing effect of the quark one-loop vacuum fluctuation on the strength of chiral transition, is somewhat larger in the QMVT model.

### v2

@@ III.3 Comparing various aspects of Columbia plots in the QMVT and RQM model for different $\bf{m_{\sigma}}$
The Fig. (7(a)) depicts, the comparison of the lines and regions in the horizontal $m_{\pi}-m_{K}$ plane (at $\mu=0$) of the three dimensional Columbia plots of the QMVT (RQM) model in the Fig. (4) (Fig. (4)) for the $m_{\sigma}=400$ MeV and Fig. (5) (Fig. (5)) for the $m_{\sigma}=530$ MeV. The Z(2) critical black solid line, in the horizontal $\mu-m_{K}$ plane of the above figures, has been replotted, by the dash dotted deep green line (solid red line ) for the $m_{\sigma}=530(400)$ MeV in the QMVT model and the dash double dotted sky blue line (solid indigo line) for the $m_{\sigma}=(530)400$ MeV in the RQM model. It is recalled that the QMVT and RQM model Z(2) critical lines for the $m_{\sigma}=400(530)$ MeV, intersect the SU(3) symmetric $m_{\pi}=m_{K}$ chiral limit line at the critical pion mass $m_{\pi}^{c}$ and terminate on the strange chiral limit line satisfying the $h_{y}=0=h_{y0}$, at the terminal pion mass $m_{\pi}^{t}$. One finds the $(m_{\pi}^{t},m_{\pi}^{c})=(98.15,78.45)$($(m_{\pi}^{t},m_{\pi}^{c})=(90.03,72.12)$) MeV in the QMVT model while in the RQM model one gets the $(m_{\pi}^{t},m_{\pi}^{c})=(168.39,134.16)$ ($(m_{\pi}^{t},m_{\pi}^{c})=(160.75,128.38)$) MeV for the $m_{\sigma}=400(530)$ MeV. The Fig. (7(a)) clearly demonstrates that the first order area underneath the crossover region of the QMVT model, is significantly smaller than the corresponding first order area lying below the crossover region of the RQM model. Table (3) shows the comparison of our present QMVT model results with those of the e-MFA:QM model FRG work Resch and RQM model work reported very recently in the Ref. vkt25 . The FRG work reports the $(m_{K}^{TCP},~m_{\pi}^{t},~m_{\pi}^{c})=(169,~110,~86)$ MeV for the $m_{\sigma}=530$ MeV, and we are finding $(m_{K}^{TCP},~m_{\pi}^{t},~m_{\pi}^{c})=(124.7,~90.03,~72.12)$ MeV in the present QMVT model work while one notes the $(m_{K}^{TCP},~m_{\pi}^{t},~m_{\pi}^{c})\equiv(235.7,~160.75,~128.38)$ MeV in the RQM model work. The critical quantities, computed in the RQM model with the large $N_{c}$ standard U(3) ChPT input set $\text{M}_{\eta}$-II when the $m_{\sigma}=530$ MeV, are slightly different as the $(m_{K}^{TCP},~m_{\pi}^{t},~m_{\pi}^{c})\equiv(239.58,~161.35,~128.04)$ MeV vkt25 . While being much smaller than their numerical values computed in the RQM model, the critical quantities $(m_{K}^{TCP},~m_{\pi}^{t},~m_{\pi}^{c})$ found in the present QMVT model work, are somewhat smaller but stand closer to the corresponding values reported in the e-MFA QM model FRG study of the Ref.Resch . The tricritical lines in the $\mu-m_{K}$ plane, show similar saturation trends for the QMVT model and the e-MFA:QM model FRG study while the saturation trend for the RQM model tricritical line gets spoiled.
The first order regions in the $\mu-m_{K}$ and $m_{\pi}-m_{K}$ planes of the QMVT model Columbia plot,  while being much smaller than the corresponding first order regions for the RQM model, are similar but moderately smaller than the extent of the first order regions computed for the e-MFA:QM model FRG study. Here, it is relevant to point out that the $\sigma$ mass, in the e-MFA:QM model FRG study, is calculated from the curvature of the effective potential that gets evaluated by the FRG flow equation. Recall that the curvature masses of mesons are used for fixing the parameters in the QMVT model and this context of incorporating quark one-loop vacuum correction in the QMVT model is analogous to that of the e-MFA:QM model FRG study in the Ref.Resch . It is important to point out that even though the fraction $\beta_{cep}$, by which the pion and kaon masses are to be reduced in the $m_{\pi}-m_{K}$ plane such that the chiral crossover ends into a second order critical end point giving the first order transition region afterwards when $\beta<\beta_{cep}$, is almost the same in both the models as $(\beta_{cep}|_{m_{\sigma}=530\text{MeV}:\text{QMVT}}=0.2028)\simeq(\beta_{cep}|_{m_{\sigma}=530\text{MeV}:\text{eMFA;QM-FRG}}\equiv 0.20)$, the extent first order transition regions in the QMVT model Columbia plot, is somewhat smaller than what is found in the e-MFA:QM model Columbia plot as the critical quantities $(m_{K}^{TCP},~m_{\pi}^{t},~m_{\pi}^{c})|_{\text{QMVT}}=(124.7,~90.03,~72.12)<(m_{K}^{TCP},~m_{\pi}^{t},~m_{\pi}^{c})|_{\text{eMFA:QM-FRG}}=(169,~110,~86)$ MeV for the $m_{\sigma}=530$ MeV. The possible reason of the above difference lies in the different methods adopted for performing chiral limit study. Recall that the $\sigma$ curvature mass $m_{\sigma}=530$ MeV does not change towards the chiral limit in the QMVT model when the ($m_{\pi},m_{K}$) dependent ChPT scaling relations for the $f_{\pi},f_{K}\text{ and }M_{\eta}$, are used for calculating the change in model parameters away from the physical point whereas in contrast, the vacuum value of the $m_{\sigma}$ decreases significantly towards the chiral limit in the e-MFA:QM model work where the $f_{\pi},M_{\eta}$ and the model parameters are kept fixed as at the physical point when the scale $\Lambda$ of the initial effective action for the FRG flow is successively increased to $\Lambda=700,1000,1100\text{ and }1143$ for compensating the successive decrease in the chiral symmetry breaking source strengths $j_{x}\text{ and }j_{y}$ by the fractions $\alpha=\sqrt{\beta}=1.0,0.17,0.04\text{ and }0$. The quark one-loop vacuum fluctuation generates quite a large smoothing effect on the strength of chiral transition in the e-MFA:QM model FRG study and this softening of the chiral transition by the fluctuation is very large (larger than that of the e-MFA:QM FRG study) in the QMVT model. Note that the full FRG QM model study where meson thermal and vacuum quantum fluctuations are also included, finds a very small $m_{\pi}^{c}\equiv 17$ MeV for which Pisarski et. al. pisarski24 has observed that this is partly because the approximation used in Ref. Resch is known to overestimate the masonic fluctuations that tend to soften the chiral transition PawlRen . Note also that the Ref. fejoHastuda has cautioned that the LPA in the FRG studies, completely neglects the wave function renormalization which might affect the final result. Recall that the quark one-loop vacuum fluctuation softens the chiral transition moderately in the RQM model because here the on-shell renormalization of parameters gives rise to stronger $U_{A}(1)$ anomaly strength $c$ together with a weaker light and strange chiral symmetry breaking strengths $h_{x}\text{ and }h_{y}$. Hence one gets significantly large first order regions in the RQM model Columbia plots vkt25 .
@@
Figure 10: Left panel (a) (Right panel (b)) shows the comparison of lines and regions in the vertical $\mu-m_{s}$, $m_{ud}=0$ (horizontal $m_{ud}-m_{s}$, $\mu=0$) planes of the three dimensional QMVT and RQM model Columbia plots for the $m_{\sigma}=400,\ 530$ MeV in Fig.(8), Fig.(9) and Fig. (8), Fig.(9). QMVT and RQM model tricritical lines for the $m_{\sigma}=400\text{ and }530$ MeV, as marked and explained in Fig.(a), start from the $m_{s}^{TCP}$ on the $m_{s}$ axis and demarcate the first order region above them on their left hand side from the second order region below them on their right hand side. QMVT and RQM model Z(2) critical lines for the $m_{\sigma}=400\text{ and }530$ MeV, as marked and explained in Fig.(b), separating the crossover transition region above them from the first order region below them, intersect the SU(3) symmetric $m_{ud}-m_{s}$ chiral limit line at the critical light quark mass $m_{ud}^{c}$ and terminate  at the terminal light quark mass $m_{ud}^{t}$.
@@ IV Summary
The consistent chiral limit has been explored in the QMVT model schafwag12 ; chatmoh1 ; vkkt13 ; Rai whose parameters get fixed by the use of curvature masses of mesons after the quark one-loop vacuum correction is included in its effective potential in the $\overline{MS}$ scheme using the extended mean field approximation (e-MFA). In the RQM model treatment of quark one-loop vacuum fluctuation, one finds the renormalized parameters by matching the counter terms in the $\overline{MS}$ scheme with those of the on-shell scheme vkkr22 ; skrvkt24 . Using the $\mathcal{O}(\frac{1}{f^{2}})$ accurate results of the infrared regularized U(3) ChPT Herpay:05 ; borasoyI ; borasoyII ; Beisert ; Becher scaling relations for the ($m_{\pi},m_{K}$) dependent expressions of the $f_{\pi},f_{K}\text{ and }M_{\eta}$, the parameter fixing for the QMVT and RQM models, have been rendered free from any ambiguity or heuristic adjustment as the $\pi,K$ meson masses are reduced to move away from the physical point for performing the chiral limit study. After generating the QMVT model Columbia plots for the $m_{\sigma}=400\text{ and }530$ MeV, these plots have been compared and contrasted with the Columbia plots generated respectively in the e-MFA:QM model FRG study Resch for the $m_{\sigma}=530$ MeV and the RQM model study for the $m_{\sigma}=400\text{ and }530$ MeV as reported very recently in the Ref.vkt25 .
@@
Very close to the chiral limit when $\beta=0.00725$, one gets huge first order gap in the QMVT and RQM model order parameters. Note that for the smaller $\beta<\beta_{cep}$, the difference in the QMVT and RQM model $T_{c}^{\chi}$ decreases from about 24 MeV for the $\beta=1.0\text{ and }0.5$ to 16.35 MeV for $\beta=0.15$ and 13.35 MeV for $\beta=0.00725$. In the exact chiral limit $f_{\pi}=f_{K}$, the $m_{\eta}=0$ and the heavy $m_{\eta^{\prime}}\equiv 825$ MeV signifies the $U_{A}(1)$ anomaly for both the QMVT model and RQM model. The spontaneous chiral symmetry breaking ( SCSB) for the $m_{\sigma}=400-800$ MeV, does not get lost in the QMVT and RQM model when the infrared regularized U(3) ChPT scaling Herpay:05 of the $m_{\pi},m_{K},M_{\eta}^{2}$ is used for the $(m_{\pi},m_{K}\rightarrow 0)$ chiral limit study. We point out that the $\sigma$ meson vacuum curvature mass stays the same as $m_{\sigma}=530$ MeV when the fraction $\beta$ is reduced in the QMVT model, whereas in the RQM model, the $\sigma$ curvature mass $m_{\sigma,c}$  is lowest at the physical point and it increases successively when one moves towards the chiral limit by reducing $\beta$. In complete contrast of the above two scenarios, the $\sigma$ mass in the e-MFA:QM model FRG study, registers a significant successive decrease towards the chiral limit from its physical point value of $m_{\sigma}=530$ MeV.
@@
The second order transition turns first order at the tricritical point $m_{K}=m_{K}^{TCP}$ ($m_{s}=m_{s}^{TCP}$) on the $m_{K}(m_{s})$ axis of the Columbia plot. The line of $Z(2)$ critical points, intersects the SU(3) symmetric $m_{\pi}=m_{K}$ ($m_{ud}=m_{s}$) line and terminates on the strange chiral limit line $h_{y0}=0(m_{s}=0)$ respectively at the $m_{\pi}^{c}$ ($m_{ud}^{c}$) and $m_{\pi}^{t}$ ($m_{ud}^{t}$) known as critical and terminal $\pi(\text{light quark})$ mass. In the $m_{\pi}-m_{K}$ and $\mu-m_{K}$ plane of the QMVT model Columbia plot for the $m_{\sigma}=400$ MeV, the computed critical quantities in the present work are $(\ m_{K}^{TCP},m_{\pi}^{t},m_{\pi}^{c})=(137.2,98.15,78.45)$ MeV which give the corresponding critical quantities in the $m_{ud}-m_{s}$ and $\mu-m_{s}$ plane as $(m_{s}^{TCP},m_{ud}^{t},m_{ud}^{c})=(8.06,2.05,1.31)$ MeV. In the RQM model $m_{\pi}-m_{K}$ and $\mu-m_{K}$ plane of the Columbia plot when the $m_{\sigma}=400$ MeV, one finds $(\ m_{K}^{TCP},m_{\pi}^{t},m_{\pi}^{c})=(242.6,168.39,134.16)$ MeV which give the $(m_{s}^{TCP},m_{ud}^{t},m_{ud}^{c})=(25.53,6.13,3.84)$ MeV in the $m_{ud}-m_{s}$ and $\mu-m_{s}$ plane. For the $m_{\sigma}=530$ MeV in the RQM model, the critical quantities reported in the Ref. vkt25 , are $(\ m_{K}^{TCP},m_{\pi}^{t},m_{\pi}^{c})=(235.7,160.75,128.38)$ MeV which correspond to the light-strange quark mass plane critical quantities as $(m_{s}^{TCP},m_{ud}^{t},m_{ud}^{c})=(24.08,5.58,3.51)$ MeV. In the present QMVT model Columbia plot calculation for the $m_{\sigma}=530$ MeV, we are finding the $(\ m_{K}^{TCP},m_{\pi}^{t},m_{\pi}^{c})=(124.7,90.03,72.12)$ MeV which give the $m_{ud}-m_{s}$ and $\mu-m_{s}$ plane critical quantities as $(m_{s}^{TCP},m_{ud}^{t},m_{ud}^{c})=(6.64,1.72,1.1)$ MeV. Note that the critical light quark mass $m_{ud}^{c}=1.1-1.31$ MeV is small in the QMVT model and it is $m_{ud}^{c}=3.51-3.84$ MeV in the RQM model whereas the recent LQCD study in Ref. zhang24 using Möbius domain wall fermions finds that $m_{ud}^{c}\leq$4 MeV.
@@
Comparing the QMVT and the e-MFA:QM FRG model studies, we note that even though the fraction $\beta_{cep}$ is almost the same in both the models as $(\beta_{cep}|_{m_{\sigma}=530\text{MeV}:\text{QMVT}}=0.2028)\simeq(\beta_{cep}|_{m_{\sigma}=530\text{MeV}:\text{eMFA;QM-FRG}}\equiv 0.20)$, the extent of first order transition regions in the QMVT model Columbia plot, is somewhat smaller than what is found in the e-MFA:QM model FRG Columbia plot as the critical quantities $(m_{K}^{TCP},~m_{\pi}^{t},~m_{\pi}^{c})|_{\text{QMVT}}=(124.7,~90.03,~72.12)<(m_{K}^{TCP},~m_{\pi}^{t},~m_{\pi}^{c})|_{\text{eMFA:QM-FRG}}=(169,~110,~86)$ MeV for the $m_{\sigma}=530$ MeV. This difference is caused by different methods of approaching the chiral limit in the two studies.
@@
Our first goal of comparing the critical quantities in the chiral limit studies of the QMVT and RQM model, has clearly established the important result that the first order regions in the $\mu-m_{K}$($\mu-m_{s}$) and $m_{\pi}-m_{K}$($m_{ud}-m_{s}$) planes of the QMVT model Columbia plot, are much smaller than the corresponding RQM model first order regions. The large first order regions in the RQM model Columbia plots are caused by the fact that the $U_{A}(1)$ anomaly strength $c$ which contains a condensate dependent part,  gets significantly enhanced when the meson self energies due to quark loops are evaluated using the meson pole masses and parameters are fixed on-sell. Note that the strength $c$, being same as in the QM model, does not change in the curvature mass based parameter fixing of the QMVT model as well as the e-MFA:QM model FRG study.The second objective of the present study led to the conclusion that the first order regions in the QMVT model Columbia plot are similar but moderately smaller than the extent of the first order regions reported in the $\mu-m_{K}$ and $m_{\pi}-m_{K}$ planes of the Columbia plot of the e-MFA:QM model FRG study for the $m_{\sigma}=530$ MeV. The above difference is caused by different methods of approaching the chiral limit in the two studies as the $f_{\pi},f_{K},M_{\eta}$ and parameters, change towards the chiral limit in the QMVT model, whereas the $f_{\pi},~M_{\eta}$ and parameters in the e-MFA:QM-FRG study are kept fixed to their physical point value and the initial effective action is successively adjusted to the higher scales $\Lambda$ as one approaches the chiral limit.
ACKNOWLEDGMENTS
I would like to express my sincere gratitude towards Suraj Kumar Rai, Akanksha Tripathi and Pooja Kumari for the careful readings of the manuscript. I am deeply indebted and very much tankful to Swatantra Kumar Tiwari for making the three dimensional Columbia plots.